A body of mass 40kg stands on a weighing machine in an accelerated lift. The reading on the scale of the weighing machine is 300N.find the magnitude and direction of acceleration.

Grade:12

1 Answers

Praneeth Mosuru
6 years ago

weight is the upward reaction force of the base we stand on=R

gravitational force is acts downward on the body=mg

an upward acceleration exists=ma

now how to form the equation of motion, 1st forget the acceleration. without it R and mg balance each other out. net force is zero

R=mg ;

mg -R=0

now there is a force present ma, so the equation becomes

mg-R=ma

the force ma is upward, so correct the direction

R-mg=ma

ma=R-mg=800–400=400 N

then a=10 m/s2.